a. What is the present value of $240,000 to be received after 30 years with a 16 percent discount rate? Use Appendix B as an approximate answer, but calculate your final answer using the formula and financial calculator methods. (Do not round intermediate calculations. Round your final answer to 2 decimal places.)
  

b. Would the present value of the funds in part a be enough to buy a $2,800 concert ticket?
  

  • No

  • Yes
Answer #1

Solution:

a) Given,

Future value = $240,000

Period (t) = 30

Discount rate (r) = 16%

Present value = Future value ( 1 ÷ (1+r)^t)

Present value = 240,000 ( 1 ÷ (1+0.16)^30)

Present value = 240,000 (1÷ 85.85)

Present value = 240,000 (0.01165)

Present value = $2,795.57

b) answer = yes

As the present value of funds in part a. is less than $2,800. So it is enough to buy a $2,800 concert ticket.